ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 201, 2014

Ahlberg, D., E. Broman, S. Griffiths and R. Morris Noise sensitivity in continuum percolation

847

Andreatta, F., A. Iovita and G. Stevens Overconvergent modular sheaves and modular forms for GL_{2/F}

299

Arhancet, C. and C. Le Merdy Dilation of Ritt operators on L^{p}- spaces

373

Benyounes, M., E. Loubeau and R. Slobodeanu Biharmonic holomorphic maps and conformally Kähler geometry

525

Bergelson, V., G. Kolesnik, M. Madritsch, Y. Son and R. Tichy Uniform distribution of prime powers and sets of recurrence and van der Corput sets in ℤ^{k}

729

Bergweiler, W. and D. A. Nicks Foundations for an iteration theory of entire quasiregular maps

147

Berkes, I. and M. Weber On series ∑ c_{k}f(kx) and Khinchin’s conjecture

593

Bourgain, J. On toral eigenfunctions and the random wave model  611

Broman, E. See D. Ahlberg Brussel, E. and E. Tengan Tame division algebras of prime period

over function fields of p-adic curves

361

Conze, J.-P. and A. Piȩkniewska On multiple ergodicity of affine cocycles over irrational rotations

543

Dranishnikov, A. and M. Levin On dimensionally exotic maps  967
Eelbode, D. The biaxial Fueter theorem  233

Falik, D. and E. Friedgut Between arrow and Gibbard– Satterthwaite; A representation theoretic approach

247

Farah, I., B. Hart and D. Sherman Model theory of operator algebras II: model theory

477

Farah, I. and S. Shelah Trivial automorphisms  701

Finis, T. and E. Lapid Relation spaces of hyperplane arrangements and modules defined by graphs of fiber zonotopes

901

Foster, T., P. Gross and S. Payne Limits of tropicalizations  835

Friedgut, E. See D. Falik González-Meneses, J. and E. Ventura Twisted conjugacy in

braid groups

455

Griffiths, S. See D. Ahlberg Gross, P. See T. Foster Hart, B. See I. Farah Iovita, A. See F. Andreatta

Kiwi, J. Puiseux series dynamics of quadratic rational maps  631

Kojman, M. Shelah’s revised GCH theorem and a question by Alon on infinite graphs colorings

585

Kolesnik, G. See V. Bergelson Kwiatkowska, A. Large conjugacy classes, projective Fraïssé limits

and the pseudo-arc

75

Lapid, E. See T. Finis Levin, M. See A. Dranishnikov

Li, W. Δ_{2} degrees without Σ_{1} induction  989

Lima, B., N. Santos and P. Sousa Translation hypersurfaces with constant scalar curvature into the Euclidean space

797

Loke, H. Y., J. Ma and U-Liang Tang Transfers of K-types on local theta lifts of characters and unitary lowest weight modules

1

Loubeau, E. See M. Benyounes Ma, J. See H. Y. Loke Madritsch, M. See V. Bergelson McCutcheon, R. and A. Windsor D sets and a Sárközy theorem

for countable fields

123

Le Merdy, C. See C. Arhancet

Montaldo, S. and I. I. Onnis Helix surfaces in the Berger sphere  949

Morris, R. See D. Ahlberg Nicks, D. A. See W. Bergweiler Onnis, I. I. See S. Montaldo Papageorgiou, N. S. and F. Papalini Seven solutions with sign

information for sublinear equations with unbounded and indefinite potential and no symmetries

761

Papalini, F. See N. S. Papageorgiou Payne, S. See T. Foster Piȩkniewska, A. See J.P. Conze

Pikhurko, O. On possible Turán densities  415
Puder, D. Primitive words, free factors and measure preservation  25
Rainer, A. Differentiable roots, eigenvalues, and eigenvectors  99

Roychowdhury, M. K. Hausdorff and upper box dimension estimate of hyperbolic recurrent sets

507

Santos, N. See B. Lima Seigal, A. L. and S. Yakovenko Local dynamics of intersections:

V. I. Arnold’s theorem revisited

813

Shelah, S. Pseudo PCF  185

Shelah, S. See I. Farah Sherman, D. See I. Farah Slobodeanu, R. See M. Benyounes Son, Y. See V. Bergelson Sousa, P. See B. Lima Stevens, G. See F. Andreatta Tang, U-Liang See H. Y. Loke Tengan, E. See E. Brussel Tichy, R. See V. Bergelson Ventura, E. See J. González-Meneses Weber, M. See I. Berkes Windsor, A. See R. McCutcheon Yakovenko, S. See A. L. Seigal